# Determinism and Asymmetry in General Relativity

### v1

@@ 5 De Dicto/De Re + Local Uniqueness
We now consider a question analogous to one posed at the end of the previous section. It concerns a connection between de dicto**/de re** determinism and the theorem of Choquet-Bruhat and Geroch (1969). Recall that $\mathscr{V}^{+}$ is the collection of four-dimensional, inextendible, vacuum, globally hyperbolic spacetimes. Which sub-collections of $\mathscr{V}^{+}$ do count as de dicto**/de re** deterministic? Because any collection $\mathscr{C}\subseteq\mathscr{H}$ that is both Heraclitus and de dicto deterministic must be de dicto** deterministic, Propositions 2 and 6 imply the following.
@@
Corollary 4. Any Heraclitus collection $\mathscr{C}\subseteq\mathscr{V}^{+}$ is de dicto** deterministic and thus de re** deterministic.
@@
As before, one naturally wonders: is there a non-empty Heraclitus collection $\mathscr{C}\subseteq\mathscr{V}^{+}$? As with the giraffe case, presumably there is. The collection of all spacetimes with trivial local symmetries is a Heraclitus collection and there is reason to think that its spacetimes are, in some sense, “generic” among the collection $\mathscr{U}$ of all spacetimes (see Sunada 1985, Proposition 1).
@@
We close by drawing attention to a strong connection between de dicto**/ de re** determinism and another (more metaphysical) way of capturing MLE determinism which requires an “intrinsic” agreement of possible worlds. This form of determinism goes by many names. It is unfortunately called “de re” by Hawthorne (2006). It is also called “Det1” by Pooley (2021) and “full” by Teitel (2019) and Halvorson et al. (2025). To avoid confusion, let us follow the latter usage: a collection of possible worlds satisfies “full determinism” if, for any worlds in the collection that “intrinsically” agree on some initial segment, the worlds “intrinsically” agree in their entirety.
@@
It would seem that in order to make sense of “intrinsic” agreement, one must restrict attention to theories in which objects are somehow named, e.g. the $\beta$ particles in Belot’s particle decay example (see Halvorson et al. 2025). Due to the possibility of symmetries, collections of general relativistic spacetimes will not have such names in general. But we emphasize here that collections of general relativistic spacetimes that satisfies de dicto**/de re** determinism are necessarily Heraclitus. This means that each point in each spacetime is effectively named since isometries (even local isometries) among Heraclitus spacetimes are unique. In Example 4, one can verify that each spacetime point can be distinguished from any other via its unique invariant curvature properties. So we see that a collection of spacetimes that is sufficiently asymmetric will have sufficiently many names to make sense of the “intrinsic” agreement of possible worlds. It follows that a collection of general relativistic spacetimes that satisfies de dicto**/de re** determinism will also satisfy full determinism (Halvorson et al. 2025). Since we know that there are collections of general relativistic spacetimes that satisfy de dicto**/de re** determinism (e.g. the one given in Example 4), we know that there are collections of general relativistic spacetimes that satisfy full determinism.
6 Conclusion
@@
We have considered several precise formulations of MLE determinism idea: if worlds agree on some initial segment, then the worlds agree entirely. We emphasize that we have not tried to capture, once and for all, a single notion of “determinism” within general relativity. In a pluralistic way, we have simply mapped out which collections of general relativistic spacetimes are deterministic relative to which definitions. We have also clarified the relationships between the definitions and posed some open questions. Here is a summary of the situation.
@@
The de dicto and de re formulations of MLE determinism have long played a central role in discussions of the hole argument (Butterfield 1989; Melia 1999; Pooley 2021). Within that context, de re* determinism was first introduced by Belot (1995). This definition introduces a uniqueness clause into the consequent of the de re form of MLE determinism: worlds must agree in only one way. Within this context, this uniqueness clause amounts to the asymmetry condition of rigidity (Geroch 1969) which has recently become a focus of attention in the hole argument literature (Halvorson and Manchak 2022; Menon and Read 2023; Dewar 2024). Unlike the de dicto and de re forms of determinism, the de re* form is strong enough to capture the intuition of Earman (1977) that Leibnizian spacetime is indeterministic. Even so, by the lights of certain haecceitist positions, de re* determinism is still too weak in the sense that Belot’s particle decay example is wrongly counted as deterministic under this form.
@@
One can certainly object to the significance of the particle decay example (Brighouse 1997). Indeed, it has been argued that one cannot hold both of the following: (i) the example counts as de re* deterministic and (ii) it makes sense to say that “$\beta_{1}$ could have been $\beta_{2}$” (Halvorson et al. 2025). But in this paper, we have explored an altogether different type of response. We have looked for even stronger versions of MLE determinism to rule out such examples. We first considered de dicto* determinism which is the analogue to re de* determinism. This definition introduces a uniqueness clause which amounts to the giraffe asymmetry condition (Barrett et al. 2023). We found that de dicto* determinism is neither stronger nor weaker than de re* determinism but that it is strong enough to rule out the globally symmetric problem cases (e.g. the particle decay example) central to the hole argument literature. Even so, it still may not be strong enough by the lights of certain haecceitist positions: some locally symmetric examples that seem to be indeterministic (e.g. the collapsing columns) are not counted as such under the de dicto* form.
@@
We then introduced a local uniqueness clause into both the de dicto and de re forms of determinism: even parts of worlds (including initial segments) must agree in only one way. This local uniqueness clause amounts to the Heraclitus asymmetry condition (Manchak and Barrett 2023). The resulting definitions of de dicto** and de re** determinism turn out to be equivalent and also seem to be sufficiently strong to rule out, not only the globally symmetric problem cases, but locally symmetric variants as well. We have shown that the satisfaction of de dicto**/re re** determinism by a collection of general relativistic spacetimes implies that each point in each spacetime is effectively named since isometries (even local isometries) among Heraclitus spacetimes are unique. This allows one to make sense of “intrinsic” agreement of possible worlds which is required by “full” determinism (Hawthorne 2006; Teitel 2019; Pooley 2021). It follows that any collection of general relativistic spacetimes that satisfies de dicto**/de re** determinism will also satisfy full determinism. We have shown that de dicto**/de re** determinism, strong as it is, is weak enough to be satisfied by some non-empty collections of general relativistic spacetimes. It is an open question whether there are large swaths of general relativistic spacetimes of physical interest that are de dicto**/de re** deterministic. Perhaps there are given that the Heraclitus condition seems to be generically satisfied (see Sunada 1985, Proposition 1).
Acknowledgments
@@
Thanks to David Malament and Neil Dewar for helpful comments on a previous draft. JOW: This material is based upon work supported by the National Science Foundation under Grant No. 2419967.

### v2

@@ 5. De Dicto/De Re + Local Uniqueness
We now consider a question analogous to one posed at the end of the previous section. It concerns a connection between de dicto**/de re** determinism and the theorem of Choquet-Bruhat and Geroch ([1969]). Recall that $\mathscr{V}^{+}$ is the collection of four-dimensional, inextendible, vacuum, globally hyperbolic spacetimes. Which sub-collections of $\mathscr{V}^{+}$ do count as de dicto**/de re** deterministic? Because any collection $\mathscr{C}\subseteq\mathscr{H}$ that is both Heraclitus and de dicto deterministic must be de dicto** deterministic, Propositions 2 and 6 imply the following:
@@
Corollary 4: Any Heraclitus collection $\mathscr{C}\subseteq\mathscr{V}^{+}$ is de dicto** deterministic and thus de re** deterministic.
@@
As before, one naturally wonders: is there a non-empty Heraclitus collection $\mathscr{C}\subseteq\mathscr{V}^{+}$? As with the giraffe case, presumably there is. The collection of all spacetimes with trivial local symmetries is a Heraclitus collection and there is reason to think that its spacetimes are, in some sense, ‘generic’ among the collection $\mathscr{U}$ of all spacetimes (see Sunada [1985], Proposition 1).
@@
We close by drawing attention to a strong connection between de dicto**/ de re** determinism and another (more metaphysical) way of capturing MLE determinism which requires an ‘intrinsic’ agreement of possible worlds. This form of determinism goes by many names. It is unfortunately called ‘de re’ by Hawthorne ([2006]). It is also called ‘Det1’ by Pooley ([2021]) and ‘full’ by Teitel ([2019]) and Halvorson et al. ([unpublished]). To avoid confusion, let us follow the latter usage: a collection of possible worlds satisfies ‘full determinism’ if, for any worlds in the collection that ‘intrinsically’ agree on some initial segment, the worlds ‘intrinsically’ agree in their entirety.
It would seem that in order to make sense of ‘intrinsic’ agreement, one must restrict attention to theories in which objects are somehow named, e.g. the $\beta$ particles in Belot’s particle decay example. This is because full determinism requires that the theory specify which object in its domain has what future properties (see Halvorson et al. ([unpublished]) for an extended discussion of this matter). Due to the possibility of symmetries, collections of general relativistic spacetimes will not have such names in general. (Here, we are following standard practice and assuming that names (whatever they may be – properties, constants, etc) must be preserved by model isomorphisms. So if all points have names, then ipso facto the model is not symmetric.)
@@
But we emphasize that collections of general relativistic spacetimes that satisfy de dicto**/ de re** determinism are necessarily Heraclitus. This means that each point in each spacetime is effectively named since isometries (even local isometries) among Heraclitus spacetimes are unique. In Example 4, one can verify that each spacetime point can be distinguished from any other via its unique invariant curvature properties. So we see that a collection of spacetimes that is sufficiently asymmetric will have sufficiently many names to make sense of the ‘intrinsic’ agreement of possible worlds. It follows that a collection of general relativistic spacetimes that satisfies de dicto**/de re** determinism will also satisfy full determinism (Halvorson et al. [2025]). Since we know that there are collections of general relativistic spacetimes that satisfy de dicto**/de re** determinism (e.g. the one given in Example 4), we know that there are collections of general relativistic spacetimes that satisfy full determinism.
6. Conclusion
@@
We have considered several precise formulations of MLE determinism idea: if worlds agree on some initial segment, then the worlds agree entirely. We emphasize that we have not tried to capture, once and for all, a single notion of ‘determinism’ within general relativity. In a pluralistic way, we have simply mapped out which collections of general relativistic spacetimes are deterministic relative to which definitions. We have also clarified the relationships between the definitions and posed some open questions. Here is a summary of the situation.
@@
The de dicto and de re formulations of MLE determinism have long played a central role in discussions of the hole argument (Butterfield [1989]; Melia [1999]; Pooley [2021]). Within that context, de re* determinism was first introduced by Belot ([1995]). This definition introduces a uniqueness clause into the consequent of the de re form of MLE determinism: worlds must agree in only one way. Within this context, this uniqueness clause amounts to the asymmetry condition of rigidity (Geroch [1969]) which has recently become a focus of attention in the hole argument literature (Menon and Read [2023]; Dewar [2025]; Halvorson and Manchak [2025]). We have seen how a generalized version of de re* determinism is a sufficient condition on spacetime theories to ensure that the identity map is not an isomorphism between models related by a non-trivial hole diffeomorphism (cf. Weatherall [2018]). The same cannot be said for generalized versions of de re or de dicto determinism. Moreover, unlike the de dicto and de re forms of determinism, the de re* form is strong enough to capture the intuition of Earman ([1977]) that Leibnizian spacetime is indeterministic. Even so, by the lights of certain haecceitist positions, de re* determinism is still too weak in the sense that Belot’s particle decay example is wrongly counted as deterministic under this form.
@@
One can object to the significance of the particle decay example (Brighouse [1997]). Indeed, it has been argued that one cannot hold both of the following: (i) the example counts as de re* deterministic and (ii) it makes sense to say that $\beta_{1}$ could have been $\beta_{2}$’ (Halvorson et al. [unpublished]). But in this paper, we have explored an altogether different type of response. We have looked for even stronger versions of MLE determinism to rule out such examples. We first considered de dicto* determinism which is the analogue to re de* determinism. This definition introduces a uniqueness clause which amounts to the giraffe asymmetry condition (Barrett et al. [2023]). We found that de dicto* determinism is neither stronger nor weaker than de re* determinism but that it is strong enough to rule out the globally symmetric problem cases (e.g. the particle decay example) central to the hole argument literature. Even so, it still may not be strong enough by the lights of certain haecceitist positions: some locally symmetric examples that seem to be indeterministic (e.g. the collapsing columns) are not counted as such under the de dicto* form.
@@
We then introduced a local uniqueness clause into both the de dicto and de re forms of determinism: even parts of worlds (including initial segments) must agree in only one way. This local uniqueness clause amounts to the Heraclitus asymmetry condition (Manchak and Barrett [forthcoming]). The resulting definitions of de dicto** and de re** determinism turn out to be equivalent and also seem to be sufficiently strong to rule out, not only the globally symmetric problem cases, but locally symmetric variants as well. We have shown that the satisfaction of de dicto**/re re** determinism by a collection of general relativistic spacetimes implies that each point in each spacetime is effectively named since isometries (even local isometries) among Heraclitus spacetimes are unique. This allows one to make sense of ‘intrinsic’ agreement of possible worlds which is required by ‘full’ determinism (Hawthorne [2006]; Teitel [2019]; Pooley [2021]). It follows that any collection of general relativistic spacetimes that satisfies de dicto**/de re** determinism will also satisfy full determinism. We have shown that de dicto**/de re** determinism, strong as it is, is weak enough to be satisfied by some non-empty collections of general relativistic spacetimes. It is an open question whether there are significant swathes of general relativistic spacetimes of physical interest that are de dicto**/de re** deterministic. Perhaps there are given that the Heraclitus condition seems to be generically satisfied (see Sunada [1985], Proposition 1).
Acknowledgements
@@
Thanks to Jeremy Butterfield, Neil Dewar, David Malament, Oliver Pooley, and two anonymous referees for helpful comments on previous drafts. JOW: This material is based upon work supported by the National Science Foundation under Grant No. 2419967.
JB Manchak
Department of Logic and Philosophy of Science
University of California, Irvine
Irvine, CA, USA
jmanchak@uci.edu
TW Barrett
Department of Philosophy
University of California, Santa Barbara
Santa Barbara, CA, USA
tbarrett@philosophy.ucsb.edu
HP Halvorson
Department of Philosophy
Princeton University
Princeton, NJ, USA
hhalvors@princeton.edu
JO Weatherall
Department of Logic and Philosophy of Science
University of California, Irvine
Irvine, CA, USA
weatherj@uci.edu